OpenStudy (coconut):

What is the solution to the equation log4 x + log4 2 = 3

OpenStudy (anonymous):

Write it as one log to get: log4(2x) = 3. Expo both sides by 4 to get: 2x = 4^3 = 64 x = 32.

OpenStudy (inkyvoyd):

hint: Log a+log b=log (ab) given that the two ar ein the same base.

OpenStudy (inkyvoyd):

ahhhg

OpenStudy (coconut):